OpenStudy (anonymous):

find the slope of the line 2x- 3y = 6

jimthompson5910 (jim_thompson5910):

2x- 3y = 6 - 3y = 6-2x - 3y = -2x+6 y = 2/3x-2 So the slope is -2/3

jimthompson5910 (jim_thompson5910):

oops meant to say slope is 2/3

OpenStudy (aroub):

-3y=-2x+6 y=2/3x-2 slope=2/3

OpenStudy (anonymous):

y=2/3x+6